# Hessian spectrum at the global minimum of the spherical pure-like mixed $p$-spin glasses

### v2

@@ 2 Pure $p$-spin model
We endow the unit sphere $S^{N-1}$ with the standard Riemannian structure, induced by the Euclidean Riemannian metric on $\mathbb{R}^{N}$. We fix an orthonormal frame field $E=\left(E_{i}\right)_{i=1}^{N-1}$ on $S^{N-1}$, which means that $\left(E_{i}(\bm{\sigma})\right)_{i=1}^{N-1}$ is a local orthonormal basis of $T_{\bm{\sigma}}S^{N-1}$ for any $\bm{\sigma}\in S^{N-1}$. We then define the spherical gradient and Hessian of $h_{N}(\bm{\sigma})$ as
Furthermore, let $\partial_{r}h_{N}(\bm{\sigma})$ denote the radial derivative of $h_{N}(\bm{\sigma})$ at $\bm{\sigma}\in S^{N-1}$, which can be understood as the directional derivative of $h_{N}(\bm{\sigma})$ in Euclidean space in the direction $\bm{\sigma}$. For the relationship between $h_{N}(\bm{\sigma})$ and its derivatives, we refer the reader to [Be22, Lemma 3.2].
@@
where $\varphi_{\nabla_{\mathrm{sp}}h_{N}(\sigma)}(0)$ is the density of $\nabla_{\mathrm{sp}}h_{N}(\sigma)$ at 0, $\omega_{N}=2\pi^{N/2}/\Gamma(N/2)$ is the surface area of the $(N-1)$-dimensional unit sphere. For the pure $p$-spin model, since the covariance matrix of $h_{N}(\bm{\sigma})$ and $\partial_{r}h_{N}(\bm{\sigma})$ is degenerate, and thus we trivially have
@@
$\begin{split}\mathbb{E}\operatorname{Crt}_{N}(B,D)=C_{p,N}\int_{D}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{E}\left(\frac{y}{p}\right)\right\}\mathrm{d}y,\end{split}$ (2.6)
@@
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-c_{p}+r_{p}\right|>\delta\right\}.$
@@
$A_{N}^{p}:=\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right.$
@@
Since $c_{p}=pE_{0}(p)>2\sqrt{p(p-1)}$ by (2.11), we may choose $\varepsilon>0$ to be sufficiently small such that for all $N$ large enough and any $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$
@@
Note that for $|y+c_{p}|\leq p\varepsilon$ and a function $h$ with $\|h\|_{L}\leq 1$, for $N$ large enough
@@
$\leq\frac{\sqrt{Np}\varepsilon}{\sqrt{(N-1)p-1}}\leq 2\varepsilon.$ (2.15)
@@
$d\left(\sigma_{\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},\frac{2\sqrt{N}}{\sqrt{N-1}}},\sigma_{c_{p}^{\prime},2}\right)<\frac{\delta}{3\sqrt{p(p-1)}}.$
By the triangle inequality, for any $N$ large enough and $\varepsilon\leq\frac{\delta}{6\sqrt{p(p-1)}}$, if the event $A_{N}^{p}$ occurs, we deduce
@@
By symmetry, property (2.14) allows us to use the upper bound for the second moment of the determinate of a shifted GOE [Be22, Lemma 6.4] (note the different scaling for the definition of standard GOE matrix). This implies that for large enough $N$ and small enough $\varepsilon$, it holds for any $\gamma>0$ and all $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$ that
@@
Since the empirical spectral measures of GOE matrices satisfy an LDP with speed $N^{2}$ [benarous1997large], there exists a constant $C>0$ such that for $N$ large enough
@@
$\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}\leq-\eta.$ (2.20)
@@
The proof of this lemma follows the similar strategy as in Lemma 2.1. We need to address equality (2) with the event $A_{N}^{p}$ replaced by the event
$B_{N}^{p}:=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right\}.$
For $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$, once event $B_{N}^{p}$ occurs, for any $\delta>0$, we choose $\varepsilon>0$ to be sufficiently small and $N$ sufficiently large such that
@@
The LDP of the smallest eigenvalues of GOE matrices [BDG01] implies that for any $\delta>0$, there exists a $\zeta>0$ such that for $N$ large enough
@@
which yields the desired result (2.20) for any $\eta<\frac{\zeta}{2}$ by the logarithmic asymptotics formula (2.7), the fact $R(-c_{p})=0$ and the continuity of $R(y)$. ∎
@@
It is clearly that for any $\delta>0$ and $\varepsilon>0$
@@
Since $\frac{1}{N}H_{N}(\bm{\sigma}^{*})$ converges to $-E_{0}(p)$ almost surely as $N$ tends to $\infty$, the second term on the right hand side of (2.23) goes to 00 as $N$ tends to $\infty$ for any $\varepsilon>0$. If the event in the first term on the right hand side of (2.23) occurs, there exists at least one critical points $\bm{\sigma}\in S^{N-1}(\sqrt{N})$ with the restriction
@@
By the different scaling on the domain (2.1), one has for any $\bm{\sigma}\in S^{N-1}$,
@@
which completes the proof of (1.5). The argument above also can be applied to (1.6), since by Markov’s inequality, Lemma 2.2 directly gives $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N,p}^{\varepsilon,\delta}\geq 1)=0$. ∎
3 Pure-like Mixture models
@@
From (3), conditioning on $\frac{h_{N}(\bm{\sigma})}{\sqrt{N}}=x$ and $\frac{\partial_{r}h_{N}(\bm{\sigma})}{\sqrt{N}}=y$, we immediately arrive at
@@
$F(x,y)$ $:=\frac{1}{2}+\frac{1}{2}\log\left(\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}\right)-\frac{1}{2}(x,y)\Sigma^{-1}(x,y)^{\top}+\Psi_{*}\left(\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right)$
@@
Here we emphasize that for the $pure$-$like$ mixture models, $E_{0}$ is equal to $E_{0}(\xi)$ mentioned in the introduction, as claimed in footnote $2$ of [BSZ20]. The authors in [BSZ20, Theorem 1.1] proved that the sequence of ground state energy $GS_{N}$ converges to $-E_{0}$ almost surely under Assumption I, this is,
@@
We introduce the important threshold of energy level
@@
From [BSZ20, Lemma 5.2], due to $-E_{0}<-E_{\infty}$, we know that if $\xi(x)$ belongs to the $pure$-$like$ mixture class, the solution $-y_{0}$ in (3.7) is unique and
@@
Later, we will prove that for the $pure$-$like$ mixture case, $y_{0}$ determines the center of the limiting empirical spectral measures of the Hessian at global minimum $\bm{\sigma}^{*}$.
@@
We establish some properties of $F(x,y)$ when $x$ in a small neighborhood of $-E_{0}$, which is crucial for the subsequent proof.
@@
If $\xi(x)$ is a pure-like mixture, then for sufficiently small $\varepsilon>0$, there exists some constant $\alpha>0$ such that
@@
and there exists some constant $\beta>0$ such that
@@
which is a decreasing function on $(-\infty,0]$ with $\partial F/\partial y(-E_{0},-y_{0})=0$. For $y<-2\sqrt{\xi^{\prime\prime}(1)}$,
@@
where $c=\frac{\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}}{2\xi^{\prime\prime}(1)(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}>0$. Assume that $\varepsilon<(y_{0}-2\sqrt{\xi^{\prime\prime}(1)})^{2}$, then
@@
Since $\partial F/\partial y(-E_{0},\cdot)$ is decreasing in $(-\infty,0]$, for $\alpha>1$, we have
@@
Note that the function $\Psi_{*}$ is Lipschitz and denote by $L_{\Psi_{*}}$ the Lipschitz constant. A little algebra yields that the claim (3.11) holds for sufficiently small $\varepsilon$ with
@@
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-y_{0}+2\sqrt{\xi^{\prime\prime}}\right|>\delta\right\}.$
@@
By the property (3.10) in Lemma 3.1, there exists a sufficiently small $\varepsilon$ such that the supremum of the right hand side of above inequality is strictly smaller than 00. We then directly deduce for sufficiently small $\varepsilon$
@@
It remains to consider $J_{N}^{1}$. Note that $-y_{0}<-2\sqrt{\xi^{\prime\prime}(1)}$ by (3.9). The fact that $\alpha\sqrt{\varepsilon}$ tends to 00 as $\varepsilon$ goes to 00 allows us to use the similar arguments as those in (2) and (2). We omit some tedious but straightforward steps to get for any $\varepsilon$ small enough and $N$ large enough
@@
Using the Cauchy–Schwarz inequality and replacing $p(p-1)$ by $\xi^{\prime\prime}$ in estimates (2) and (2.19), there exists some constant $C>0$ such that for any $\varepsilon>0$ sufficiently small and $N$ large enough
@@
$\lim_{N\rightarrow\infty}\mathbb{E}\mathcal{B}_{N,\xi}^{\varepsilon,\delta}=0.$ (3.18)
@@
To prove this lemma, it suffices to prove that the equality (3) converges to 00 as $N$ tends to $\infty$ with the event $A_{N}^{\xi}$ replaced by the event
$B_{N}^{\xi}=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+y_{0})}{\sqrt{(N-1)\xi^{\prime\prime}}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right\}.$
Following the same partition with respect to variable $y$ as in (3), we denote the three parts by $I_{N}^{1}$, $I_{N}^{2}$ and $I_{N}^{3}$, respectively. For sufficiently small $\varepsilon$, by using the same argument as in (3.17), we have
@@
For $I_{N}^{1}$, since $y\in(y_{0}-\alpha\sqrt{\varepsilon},y_{0}+\alpha\sqrt{\varepsilon})$, once the event $B_{N}^{\xi}$ occurs, for any $\delta>0$, we take $\varepsilon>0$ to be sufficiently small and $N$ sufficiently large such that
@@
Then we deduce for any $\varepsilon$ small enough and $N$ large enough
@@
Combining the Cauchy–Schwarz inequality, second moment upper bound (2) and the LDP for smallest eigenvalues of GOE matrices (2.22), the logarithmic asymptotics formula (3.6) implies that there exists a constant $C>0$ such that for all $\varepsilon>0$ sufficiently small and $N$ large enough
$I_{N}^{1}\leq\exp\{-CN\}\exp\left\{\sup_{\begin{subarray}{c}x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)\\ y\in(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon})\end{subarray}}NF(x,y)+o(N)\right\}.$
By the key observation (3.11) in Lemma 3.1, for sufficiently small $\varepsilon$, it follows that $\lim_{N\rightarrow\infty}I_{N}^{1}=0$, which together with (3.19) finishes the proof of this lemma. ∎

### v3

@@ 2 Pure $p$-spin model
[…]
@@
$\begin{split}\mathbb{E}\operatorname{Crt}_{N}(B,D)=C_{p,N}\int_{D}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{B}\left(\frac{y}{p}\right)\right\}\mathrm{d}y,\end{split}$ (2.6)
@@
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-c_{p}+r_{p}\right|>\delta\right\},$
$\mathcal{C}_{N,p}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\right.$
$\left|\lambda_{\max}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-c_{p}-r_{p}\right|>\delta\right\}.$
@@
$A_{N}^{p}=\left\{L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right.$
@@
Since $c_{p}=pE_{0}(p)>2\sqrt{p(p-1)}$ by (2.11), we may choose $\varepsilon>0$ to be sufficiently small such that for all $N$ large enough and any $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$, we have
@@
Note that for $|y+c_{p}|\leq p\varepsilon$ and a function $h$ with $\|h\|_{L}\leq 1$, for any $N$ large enough
@@
$\leq\frac{\sqrt{Np}\varepsilon}{\sqrt{(N-1)(p-1)}}\leq 2\varepsilon.$ (2.15)
@@
$d\left(\sigma_{\frac{\sqrt{N}c_{p}}{\sqrt{(N-1)p(p-1)}},\frac{2\sqrt{N}}{\sqrt{N-1}}},\sigma_{c_{p}^{\prime},2}\right)\xrightarrow{N\rightarrow\infty}0.$
By the triangle inequality, for any $N$ large enough and $\varepsilon<\frac{\delta}{6\sqrt{p(p-1)}}$, if the event $A_{N}^{p}$ occurs, we deduce
@@
By symmetry, the property (2.14) allows us to apply the upper bound for the second moment of the determinant of a shifted GOE matrix [Be22, Lemma 6.4] (note the different scaling used in the definition of the standard GOE matrix). This implies that for large enough $N$ and small enough $\varepsilon$, it holds for any $\gamma>0$ and all $y\in(-c_{p}-p\varepsilon,-c_{p}+p\varepsilon)$ that
@@
Since the empirical spectral measures of the GOE matrices satisfy an LDP with speed $N^{2}$ [benarous1997large], there exists a constant $C>0$ such that for $N$ large enough
@@
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}\leq-\eta$ (2.20)
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathcal{C}_{N,p}^{\varepsilon,\delta}\leq-\eta.$ (2.21)
@@
The proof of this lemma follows a similar strategy as in Lemma 2.1. We need to address equality (2) with the event $A_{N}^{p}$ replaced by the event
$B_{N}^{p}=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+c_{p})}{\sqrt{(N-1)p(p-1)}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)p(p-1)}}\right\}.$
For any fixed $\delta>0$, there exists a $\varepsilon>0$ sufficiently small such that for any $N$ sufficiently large, the event $B_{N}^{p}$ implies the event
@@
The LDP for the smallest eigenvalues of the GOE matrices [BDG01] implies that for any $\delta>0$, there exists a $\zeta>0$ such that for all $N$ large enough
@@
which yields the desired result (2.20) for any $\eta<\frac{\zeta}{2}$ by the logarithmic asymptotics formula (2.7), the fact $R(-c_{p})=0$ and the continuity of $R(y)$. The proof for (2.21) is similar, since the LDP for the largest eigenvalues of the GOE matrices [BDG01] states that, for any $\delta>0$, there exists $\zeta>0$ such that for $N$ sufficiently large,
$\mathbb{P}\left(\left|\lambda_{\max}\left(\mathrm{GOE}_{N}\right)-2\right|>\frac{\delta}{2\sqrt{p(p-1)}}\right)\leq e^{-\zeta N}.$
@@
It is clear that for any $\delta>0$ and $\varepsilon>0$
@@
Since $\frac{1}{N}H_{N}(\bm{\sigma}^{*})$ converges to $-E_{0}(p)$ almost surely as $N$ tends to infinity, the second term on the right-hand side of (2.24) goes to zero as $N$ tends to infinity for any $\varepsilon>0$. If the event in the first term on the right-hand side of (2.24) occurs, there exists at least one critical points $\bm{\sigma}\in S^{N-1}(\sqrt{N})$ with the restriction
@@
Due to the different scaling on the domain (2.1), one has for any $\bm{\sigma}\in S^{N-1}$,
@@
which completes the proof of (1.7). The argument above can also be applied to obtain (1.8) and (1.9), since by Markov’s inequality, Lemma 2.2 directly gives $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N,p}^{\varepsilon,\delta}\geq 1)=0$ and $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{C}_{N,p}^{\varepsilon,\delta}\geq 1)=0$. ∎
3 Pure-like Mixture model
@@
From (2.3), conditioning on $\frac{h_{N}(\bm{\sigma})}{\sqrt{N}}=x$ and $\frac{\partial_{r}h_{N}(\bm{\sigma})}{\sqrt{N}}=y$, we immediately arrive at
@@
$F(x,y)$ $=\frac{1}{2}+\frac{1}{2}\log\left(\frac{\xi^{\prime\prime}(1)}{\xi^{\prime}(1)}\right)-\frac{1}{2}(x,y)\Sigma^{-1}(x,y)^{\top}+\Psi_{*}\left(\frac{y}{\sqrt{\xi^{\prime\prime}(1)}}\right)$
@@
Here we emphasize that for the $pure$-$like$ mixture model, $E_{0}$ is equal to $E_{0}(\xi)$ mentioned in the introduction section, as claimed in Footnote 222 of [BSZ20]. The authors [BSZ20, Theorem 1.1] proved that the sequence of ground state energy $GS_{N}$ converges to $-E_{0}$ almost surely under Assumption I, that is,
@@
We introduce an important threshold of energy level
@@
The value $E_{\infty}$ is related to $G\left(\xi^{\prime},\xi^{\prime\prime}\right)$ defined in (1.5) by $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=\sup_{y\in\mathbb{R}}F(-E_{\infty},y)$ [ABA13, (4.1)]. From [BSZ20, Lemmas 5.1], due to $-E_{0}<-E_{\infty}$, we know that if $\xi(x)$ belongs to the $pure$-$like$ mixture class, the solution $-y_{0}$ in (3.7) is unique and
@@
Later, we will prove that for the $pure$-$like$ mixture case, $y_{0}$ determines the center of the limiting empirical spectral measure of the Hessian at global minimum $\bm{\sigma}^{*}$.
@@
We establish some properties of $F(x,y)$ when $x$ is in a small neighborhood of $-E_{0}$, which is crucial for the subsequent proof. Morally speaking, the following lemma implies that the $\mathrm{argmax}$ function defined in (3.7), is continuous with respect to $x$ when $x$ lies in a small neighborhood of $-E_{0}$.
@@
If $\xi(x)$ is a pure-like mixture, then there exists some constant $\alpha>0$ such that for sufficiently small $\varepsilon>0$,
@@
and there exists some constant $\beta>0$ such that for sufficiently small $\varepsilon>0$,
@@
which is a decreasing function on (,0]($\infty$,0]( - ∞ , 0 ] with $\partial F/\partial y(-E_{0},-y_{0})=0$. For $y<-2\sqrt{\xi^{\prime\prime}(1)}$,
@@
where $c=\frac{\xi^{\prime\prime}(1)-\xi^{\prime}(1)+\xi^{\prime}(1)^{2}}{2\xi^{\prime\prime}(1)(\xi^{\prime\prime}(1)+\xi^{\prime}(1)-\xi^{\prime}(1)^{2})}>0$. For any $\varepsilon<(y_{0}-2\sqrt{\xi^{\prime\prime}(1)})^{2}$, it follows that
@@
Since $\partial F/\partial y(-E_{0},\cdot)$ is decreasing on (,0]($\infty$,0]( - ∞ , 0 ], for $\alpha>1$, we have
@@
Note that the function $\Psi_{*}$ is Lipschitz and denote by $L_{\Psi_{*}}$ the Lipschitz constant. A little algebra yields that the claim (3.11) holds for sufficiently small $\varepsilon>0$ with
@@
$\left|\lambda_{\min}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-y_{0}+2\sqrt{\xi^{\prime\prime}}\right|>\delta\right\},$
$\mathcal{C}_{N,\xi}^{\varepsilon,\delta}=\#\left\{\bm{\sigma}\in S^{N-1}\colon\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0,\;\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon),\right.$
$\left|\lambda_{\max}\left(\nabla^{2}_{\mathrm{sp}}h_{N}(\bm{\sigma})\right)-y_{0}-2\sqrt{\xi^{\prime\prime}}\right|>\delta\right\}.$ (3.14)
@@
By property (3.10) in Lemma 3.1, there exists a sufficiently small $\varepsilon>0$ such that the supremum on the right-hand side of the above inequality is strictly smaller than zero. We then directly deduce for sufficiently small $\varepsilon>0$
@@
It remains to consider $J_{N}^{1}$. Note that $-y_{0}<-2\sqrt{\xi^{\prime\prime}(1)}$ by (3.9). The fact that $\alpha\sqrt{\varepsilon}$ tends to zero as $\varepsilon$ goes to zero allows us to use arguments similar to those in (2) and (2). We omit some tedious but straightforward steps to obtain that for any $\varepsilon>0$ small enough and $N$ large enough,
@@
Using the Cauchy–Schwarz inequality and replacing the factor $p(p-1)$ by $\xi^{\prime\prime}$ in estimates (2) and (2.19), there exists a constant $C>0$ such that for any $\varepsilon>0$ sufficiently small and $N$ large enough
@@
$\lim_{N\rightarrow\infty}\mathbb{E}\mathcal{B}_{N,\xi}^{\varepsilon,\delta}=0,$ (3.19)
$\lim_{N\rightarrow\infty}\mathbb{E}\mathcal{C}_{N,\xi}^{\varepsilon,\delta}=0.$ (3.20)
@@
To prove this lemma, it suffices to prove that the equality (3) converges to zero as $N$ tends to infinity with the event $A_{N}^{\xi}$ replaced by the event
$B_{N}^{\xi}=\left\{\left|\lambda_{\min}\left(\mathrm{GOE}_{N-1}\right)-\frac{\sqrt{N}(y+y_{0})}{\sqrt{(N-1)\xi^{\prime\prime}}}+\frac{2\sqrt{N}}{\sqrt{N-1}}\right|>\frac{\sqrt{N}\delta}{\sqrt{(N-1)\xi^{\prime\prime}}}\right\}.$
Following the same partition with respect to variable $y$ as in (3), we denote the three parts by $I_{N}^{1}$, $I_{N}^{2}$ and $I_{N}^{3}$, respectively. For sufficiently small $\varepsilon>0$, using the same argument as in (3.18), we have
@@
For $I_{N}^{1}$, since $y\in(y_{0}-\alpha\sqrt{\varepsilon},y_{0}+\alpha\sqrt{\varepsilon})$, for any fixed $\delta>0$, there exists a $\varepsilon>0$ sufficiently small such that for all $N$ sufficiently large, the event $B_{N}^{\xi}$ implies the event
@@
Then we deduce for any $\varepsilon>0$ small enough and $N$ large enough
@@
Combining the Cauchy–Schwarz inequality, second moment upper bound (2) and the LDP for the smallest eigenvalues of GOE matrices (2.23), the logarithmic asymptotics formula (3.6) implies that there exists a constant $C>0$ such that for all $\varepsilon>0$ sufficiently small and $N$ large enough
$I_{N}^{1}\leq\exp\{-CN\}\exp\left\{\sup_{\begin{subarray}{c}x\in(-E_{0}-\varepsilon,-E_{0}+\varepsilon)\\ y\in(-y_{0}-\alpha\sqrt{\varepsilon},-y_{0}+\alpha\sqrt{\varepsilon})\end{subarray}}NF(x,y)+o(N)\right\}.$ (3.22)
By the key observation (3.11) in Lemma 3.1, for sufficiently small $\varepsilon>0$, it follows that $\lim_{N\rightarrow\infty}I_{N}^{1}=0$, which together with (3.21) finishes the proof of (3.19). As before, the proof of (3.20) is a trivial extension of the method used for (3.19), based on the LDP for the largest eigenvalues of GOE matrices. ∎